OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3 This is how it's set up (X___________=_

OpenStudy (anonymous):

@perl please help

OpenStudy (plasmataco):

(X-5)^2+(Y-6)^2=3^2

OpenStudy (plasmataco):

???

OpenStudy (plasmataco):

Is it correct?

OpenStudy (anonymous):

I don't know. I don't understand how to do this type of problem

OpenStudy (loser66):

@Plasmataco Please, explain him how to get it. It is better than giving out the free answer.

OpenStudy (loser66):

@Jcameron want to know how to get it?

OpenStudy (anonymous):

Yeah because I don't understand how you got that.

OpenStudy (anonymous):

@Loser66 Yes, or else I don't learn lol

OpenStudy (loser66):

center is (h, k) , so tell me , the given center is (5,6) hence h =?? and k =??

OpenStudy (anonymous):

H= 5 K=6 ?

OpenStudy (loser66):

yup, and radius =??

OpenStudy (anonymous):

Is 3^2?

OpenStudy (loser66):

nope, just 3, they give you radius =3, how can you get 3^2

OpenStudy (loser66):

ok?

OpenStudy (anonymous):

Oh sorry, I made a mistake. I was thinking about something different.

OpenStudy (anonymous):

Okay. So, how do I put that into an equation?

OpenStudy (loser66):

\(\color{red}{h =5}\), \(\color{blue}{k=6}\), \(\color {orange}{radius =3}\) Formula of a circle is \((x -\color{red}{h})^2+(y-\color{blue}{k})^2=\color{orange}{radius^2}\)

OpenStudy (loser66):

Plug the number in, pppppplease

OpenStudy (anonymous):

(x-5)^2+(y-6)^2=3^2 ?

OpenStudy (loser66):

yyyyyyyyyyyyyup

OpenStudy (anonymous):

Thank you!

OpenStudy (loser66):

np